OpenStudy (anonymous):

Find the sum of the arithmetic sequence. -10, -7, -4, -1, 2, 5, 8 11 -70 0 -7

OpenStudy (freckles):

you can add them up since there isn't that many

OpenStudy (freckles):

unless you want to do this: \[a_{n}=-10+3n\] where your first number is given when n is zero \[\sum_{n=1}^{7}(-10+3n)\] and use some formulas to figure out the sum either way

OpenStudy (anonymous):

i got 14 lol

OpenStudy (anonymous):

@freckles am i right lol do you know the answer

OpenStudy (freckles):

I'm going to fix my sum \[\sum_{n=0}^{6}(-10+3n)\] anyways \[-10+(-7)+(-4)+(-1)+2+5+8\] \[-17+(-5)+7+8 \\ -22+15\] I got something that isn't 14 let me know if I did something wrong

OpenStudy (anonymous):

oh its -7 lol

OpenStudy (freckles):

adding is fun :p

OpenStudy (anonymous):

It's -7...add -10+-7+-4+-1=-22 then add 2+5+8=15 add -15. To 22 = -7